# DivScene: Benchmarking LVLMs for Object Navigation with Diverse Scenes and Objects

### v2

@@ Abstract
Object navigation in unknown environments is crucial for deploying embodied agents in real-world applications. While we have witnessed huge progress due to large-scale scene datasets, faster simulators, and stronger models, previous studies mainly focus on limited scene types and target objects. In this paper, we study a new task of navigating to diverse target objects in a large number of scene types. To benchmark the problem, we present a large-scale scene dataset, DivScene, which contains 4,614 scenes across 81 different types. With the dataset, we build an end-to-end embodied agent, NatVLM, by fine-tuning a Large Vision Language Model (LVLM) through imitation learning. The LVLM is trained to take previous observations from the environment and generate the next actions. We also introduce CoT explanation traces of the action prediction for better performance when tuning LVLMs. Our extensive experiments find that we can build a performant LVLM-based agent through imitation learning on the shortest paths constructed by a BFS planner without any human supervision. Our agent achieves a success rate that surpasses GPT-4o by over 20%. Meanwhile, we carry out various analyses showing the generalization ability of our agent. Our code and data are available at https://github.com/zhaowei-wang-nlp/DivScene.
Figure 1: Illustration of our NatVLM agent. We build CoT explanation traces to help the LVLM to better grasp the rationale of object navigation.
@@ 1 Introduction
Object navigation has long been a challenging embodied task, where an embodied agent is required to navigate to target objects in unseen environments (Batra et al., 2020; Anderson et al., 2018). This task is fundamental to other navigation-based embodied tasks, as navigating to a target object is the agent’s preliminary step in interacting with objects in a scene. As the last few years have witnessed huge progress with large-scale scene datasets (Chang et al., 2017) and faster simulators (Kolve et al., 2017), numerous methods have been proposed for better navigation, including reinforcement learning (RL) (Ye et al., 2021), imitation learning (IL) (Ramrakhya et al., 2022), semantic maps (Zheng et al., 2022), and others (De Vries et al., 2018; Chaplot et al., 2020a).
Albeit those methods achieve state-of-the-art performance on existing object navigation tasks, they only focus on a limited set of target objects and scenes with little variety. For example, the photorealistic dataset Matterport-3D (Chang et al., 2017) only considers 21 target object types in 90 private homes. Similarly, the simulated environment ProcTHOR (Deitke et al., 2022) only considers 16 object types in four types of residential rooms (i.e., bedroom, living room, kitchen, and bathroom). With this limited scope, models often perform poorly (Zhou et al., 2023) when facing diverse unseen objects and scenes in real-world applications due to the distribution shifting.
@@
In this paper, we study a new task of building navigational agents capable of generalizing to a wide range of unseen objects in diverse scenes. We introduce a new dataset, DivScene, that features the most comprehensive range of scene types to the best of our knowledge. Specifically, we collect 81 scene types based on MIT Scenes Dataset (Quattoni & Torralba, 2009), which are further categorized into five big groups. Then, we use LLMs to automatically compose diverse house descriptions by adding attributes to scene types, such as “a bakery with tile patterned walls.” We input these descriptions into the language-guided framework, Holodeck (Yang et al., 2024), to build houses automatically with the strong ability of GPT-4 (OpenAI, 2023). Finally, we compile 4,614 houses across 81 distinct scene types on the AI2THOR platform (Kolve et al., 2017). Since we take advantage of the object assets Objaverse (Deitke et al., 2023), our houses contain over 22K different kinds of objects, making them ideal for testing navigation to diverse objects.
With DivScene, we proceed to build an embodied agent that can navigate to diverse objects in diverse types of houses. As aforementioned, existing works, such as RL-based (Maksymets et al., 2021) and IL-based (Ramrakhya et al., 2022) methods, only focus on learning in-distribution knowledge for limited target objects and scene types. Meanwhile, recent studies (Yu et al., 2023; Zhou et al., 2023; Dorbala et al., 2023) utilize LLMs as the planning backbone to navigate to diverse objects in a zero-shot manner, leveraging LLM’s commonsense knowledge for high-level guidance. However, they still face a substantial domain gap between navigation tasks and the LLM training corpus (Lin et al., 2024). Moreover, those methods (Chen et al., 2023; Zhu et al., 2024) require a captioning model to generate textual descriptions of each scene, serving as the perception of the backbone LLMs. The captioning model might miss crucial information about objects and lead to suboptimal results, like spatial or color details.
To tackle the challenge, we propose an end-to-end object navigation agent called NatVLM (Navigational Chain-of-Thought VLM), based on the large vision language model Idefics 2 (Laurençon et al., 2024). As shown in Figure 1, the model is tuned to generate the next action given the current observation, such as the the egocentric view and the agent’s status. Here, we train the LVLM with imitation learning (Brohan et al., 2022) using shortest paths from a heuristic planner. To improve the accuracy, we also manually collect complex CoT explanation traces of each prediction (Mitra et al., 2023; Ho et al., 2023) to help the LVLM understand the underlying rationale behind object navigation. By tuning the LVLM, we eliminate the domain gap between navigation tasks and the pre-training corpus. Meanwhile, our navigation is performed in an end-to-end manner with perception, circumventing the captioning model.
Existing IL-based works trained their navigational agents on large corpora of human demonstrations (Brohan et al., 2022; Wei et al., 2023), which is incredibly expensive. In contrast, our study finds that imitating the shortest paths constructed by a heuristic planner can be an effective approach to training agents based on LVLMs. Specifically, the AI2THOR platform discretizes the environment into a grid map with a fixed step size. Then, we search for the shortest path from the agent to the target objects. In total, we collect about 23K shortest-path episodes in the houses from DivScene, forming the DivTraj dataset. The dataset contains 5,707 different kinds of target objects, significantly more than other navigation datasets.
With shortest-path episodes, we perform extensive experiments and analyses to evaluate our NatVLM agent. First, we introduce several baselines using various LLMs and VLMs (Liu et al., 2024b; Touvron et al., 2023) to establish baseline performance levels. The evaluation results show that our proposed agent can navigate to diverse objects more effectively than baselines by a large margin. We carry out thorough ablation studies to show the efficacy of CoT explanation traces in action prediction. Moreover, few-shot experiments demonstrate the robustness of our agent. Last but not least, we validate the generalization ability of our agent on two out-of-distribution datasets: ProcTHOR (Deitke et al., 2022) and iTHOR (Weihs et al., 2021).
@@ 2 Related Work
Recent breakthroughs in large vision language models (LVLMs) have shown success in several domains (Li et al., 2023b; Wang et al., 2023a; Team, 2023; Jiang et al., 2024). Our work first builds a navigational agent based on LVLMs with the following related areas of study:
Visual Navigation: Visual navigation is a critical task for building intelligent agents imitating human behavior. Conventional methods depend on an environment map but struggle with the low generalization ability in unseen houses (Yamauchi, 1997). Recent works (Anderson et al., 2018) have studied more forms of the visual navigation task, like PointNav (Wijmans et al., 2019), ImageNav (Zhu et al., 2017), ObjectNav (Chaplot et al., 2020b), RoomNav (Wu et al., 2018), and visual-language navigation (Huang et al., 2023). In this work, we study the ObjectNav task and build the first end-to-end navigational agent based on LVLMs.
Policy Learning of Object Navigation: Zhu et al. (2017) first proposed a reinforcement learning (RL) method that used ResNet (He et al., 2016) to encode images. Following them, the RL-based approaches have gained attention for tackling object navigation (Anderson et al., 2018; Batra et al., 2020; Savva et al., 2017; Wahid et al., 2021; Yang et al., 2018; Druon et al., 2020; Ehsani et al., 2021). However, existing works (Ehsani et al., 2023; Guo et al., 2018) show that RL requires extensive reward shaping and is often too slow and ineffective. Imitation learning (Pomerleau, 1988; Zhang et al., 2018) presents a compelling alternative to RL as it reframes the task as a supervised learning problem. RT-1 and RT-2 (Brohan et al., 2022; 2023) scale up the multi-task data, focusing on object manipulation. Wei et al. (2023) also propose a prompt-guided imitation learning method. However, these methods face the tremendous cost of human demonstration. While SPOC (Ehsani et al., 2023) attempts to resolve the issue, they trained a transformer-based agent from scratch, requiring a substantial amount of training data. Meanwhile, their study is still limited to four room types in ProcTHOR (Deitke et al., 2022). Compared with those works, we fine-tune pre-trained LVLM to navigate diverse scenes and construct shortest-path trajectories without expensive cost.
@@
LLMs and VLMs for Visual Navigation: With the recent advancement, LLMs and VLMs (Achiam et al., 2023; Touvron et al., 2023; Lu et al., 2024) have emerged as the foundation for solving embodied tasks (Pan et al., 2023; Majumdar et al., 2024). Researchers also explored their usage for object navigation. A few methods have employed contrastive VLMs (Radford et al., 2021; Li et al., 2022), as the visual encoders (Khandelwal et al., 2022; Majumdar et al., 2022) or object-grounding tools (Gadre et al., 2023; Dorbala et al., 2023). Yu et al. (2023) utilize LLMs as the planning backbone for object navigation in a zero-shot manner. LLMs are used as high-level planners with a captioning model to provide perception. Following them, many improvements have been proposed (Cai et al., 2024b; Chen et al., 2023; Zhou et al., 2023; Shah et al., 2023). However, a substantial domain gap exists between navigation tasks and the LLM pre-training corpus (Lin et al., 2024). Meanwhile, using LLMs for navigation might lose important visual details. In contrast, we tune our end-to-end agent from an LVLM through imitation learning without those drawbacks.
Embodied Environment: Embodied environments are the foundation of embodied research. Existing environments are typically crafted through the manual labor of 3D artists (Deitke et al., 2020; Gan et al., 2020; Li et al., 2023a; Khanna et al., 2024; Tang et al., 2023; Xia et al., 2018), hard to scale up. Many works construct scenes from 3D scans (Savva et al., 2019; Ramakrishnan et al., 2021; Szot et al., 2021), but the scenes are not interactive. Meanwhile, ProcTHOR (Deitke et al., 2022) proposes a procedural method but only focuses on four scene types. Recently, Yang et al. (2024) introduced a language-guided system using GPT-4 to automatically generate customized scenes from textual house descriptions, called Holodeck. In this work, we collect diverse house descriptions with GPT-4, which are further used in Holodeck to collect houses.
@@ 3 Task Definition
In this paper, we study the object navigation task involving an agent in an environment to find the target object belonging to a given category. The object category set is denoted by $C=\{c_{0},c_{1},\ldots,c_{m}\}$, where $m$ is the number of categories. The scenes can be described by $S=\{s_{0},s_{1},\ldots,s_{n}\}$, and $n$ is the total number of scenes. In each episode, the embodied agent is initialized at a random position $p_{i}$ with rotation $r_{i}$ in a scene $s_{i}$. Then, the embodied agent is required to perform instance-level object navigation, where the agent receives a target object within the category $c_{i}$ at the position $o_{i}$. Thus, an episode can be represented as $E_{i}=\{s_{i},p_{i},r_{i},c_{i},o_{i}\}$. At each time step $t$, the embodied agent observes the environment and predicts the next action $a_{t}$. Following previous work (Yu et al., 2023; Zhu et al., 2024), the observation comprises an RGB image of the egocentric view and the agent status (i.e., its position and rotation).
Our new houses are all collected on the AI2THOR platform (Kolve et al., 2017), and thus the action space, signified as $A$, covers four actions for navigation: MoveAhead, RotateRight, RotateRight, and Done. We adopt the default settings of the AI2THOR platform. The MoveAhead action moves the agent 25 centimeters, and RotateLeft and RotateRight actions rotate the agent 90 degrees. Done is used when the agent believes it has finished the navigation task. When the agent takes the Done action or reaches the max action limit, the episode is considered successful if the distance to the target object is less than 1.5 meters. According to this setting, an environment can be formulated as a 0.25 $\times$ 0.25-meter grid map of all reachable positions.
Figure 2: Data collection process. On the left, we show the process of collecting the scene dataset. We prompt GPT-4 to collect textual house descriptions and use Holodeck to build houses automatically. On the right, we show an episode built in the house. We use BFS to find the shortest path from the initial position to the target object. Then, actions and observations are collected.
4 Data Collection for DivScene and DivTraj
Existing object navigation studies only focus on limited types of scenes and objects. To fill this gap, we first curate a large-scale scene dataset DivScene, featuring 81 scene types. We show the details in Figure 2. Then, 23K episodes with diverse target objects are sampled using a BFS-based planner, forming the DivTraj dataset.
@@ 4.1 Scene Collection for DivScene
We adopt the Holodeck (Yang et al., 2024) framework to build scenes automatically, easing human labor. Holodeck takes textual house descriptions as input and uses GPT-4 to decide the layout, styles, and object selections. To collect diverse houses, we first manually compile 81 scene types across five big groups by supplementing the MIT scene dataset (Quattoni & Torralba, 2009), like music studio and home office. We present a few houses in Figure 3 and more examples in Appendix E.
@@
Then, we build textual house descriptions based on randomly chosen scene types by adding house attributes. We consider 12 house attributes, such as room style, users of the room, etc. We randomly sample 1-3 attributes and prompt GPT-4 to assign specific values to them. A house description is then written by GPT-4, given the scene type and attribute values. Here, we prompt GPT-4 under the in-context learning setting with the standard instruction-then-exemplar prompt (West et al., 2021):
<INSTRUCTION>
<EX1-IN-TYPE><EX1-IN-ATTR> <EX1-OUT>
<EXN-IN-TYPE><EXN-IN-ATTR> <EXN-OUT>
<EXN+1-IN-TYPE><EXN+1-IN-ATTR>
where <INSTRUCTION> describes the task of writing house descriptions. <EXi-IN-TYPE> and <EXi-IN-ATTR> are the selected scene type and attributes, with output <EXi-OUT> being an exemplar description. Concrete prompts are shown in Section A.2, and we provide GPT-4 with $N=5$ exemplars to generate the description for the final test example. Also, we leave the full details of scene types and attributes in Section A.1
To encourage diversity, a new description is included only when its ROUGE-L similarity (Lin, 2004) with any existing description is less than 0.8. Invalid generations are also identified and filtered out based on heuristics (e.g., invalid output format). We show more details in Section A.3. After we collect abundant scene descriptions, we input each of them into the Holodeck framework and build a dataset of 4,614 scenes named DivScene.
Figure 3: Examples of houses in DivScene with different scene types.
4.2 Episode Collection For DivTraj
@@
We build a BFS-based planner and design a pipeline to produce expert trajectories of shortest paths. Our method has three steps to sample each episode. First, we sample an initial position and target object. Then, we use our planner to find the shortest path between them. Finally, we obtain the action sequence and corresponding environment observations. We show an example in Figure 2.
@@
As discussed in Section 3, houses in AI2THOR are discretized as grid maps of the fixed step size. We randomly sample an initial position from the grip map for the agent. Then, a target object is randomly sampled from all available objects in the environment. To encourage diversity, objects of the same type as those sampled in previous episodes are removed from the pool when sampling for new episodes in the same house. We also impose that the target objects are between 0.3 m and 2.0 m in height to ensure they are observable to the agent.
@@
Second, we build a planner capable of navigating multi-room cluttered environments with various obstacles. The planner stems from the ground truth information available in AI2-THOR, like reachable positions and objects’ coordinates. We use Breadth-First Search (BFS) on the grid map to find the shortest path from the initial position to the target object. We show more details of BFS in Section A.4. We stress that ground truth information is not available to agents at inference time and is only used to produce expert episodes for training.
@@
With the shortest path, we then derive the sequence of actions needed to achieve navigation. Basically, between two adjacent positions in the shortest path, we add a MoveAhead action if the agent’s rotation remains unaltered. Otherwise, we first use RotateRight or RotateLeft to adjust the orientation and then add a MoveAhead action (more details in Section A.5). Thus, we obtain an action sequence that can steer the agent to the target object. We execute all actions in AI2THOR and collect observations at each step, including the agent’s status and egocentric images.
4.3 Statistics of Both Datasets

### v3

# DivScene: Towards Open-Vocabulary Object Navigation with Large Vision Language Models in Diverse Scenes

@@ Abstract
Large Vision-Language Models (LVLMs) have achieved significant progress in tasks like visual question answering and document understanding. However, their potential to comprehend embodied environments and navigate within them remains underexplored. In this work, we first study the challenge of open-vocabulary object navigation by introducing DivScene, a large-scale dataset with 4,614 houses across 81 scene types and 5,707 kinds of target objects. Our dataset provides a much greater diversity of target objects and scene types than existing datasets, enabling a comprehensive task evaluation. We evaluated various methods with LVLMs and LLMs on our dataset and found that current models still fall short of open-vocab object navigation ability. Then, we fine-tuned LVLMs [^1] Our code and data are available at https://github.com/zhaowei-wang-nlp/DivScene. to predict the next action with CoT explanations. We observe that LVLM’s navigation ability can be improved substantially with only BFS-generated shortest paths without any human supervision, surpassing GPT-4o by over 20% in success rates.
DivScene: Towards Open-Vocabulary Object Navigation with Large Vision Language Models in Diverse Scenes
Zhaowei Wang1, Hongming Zhang2, Tianqing Fang1,2, Ye Tian3, Yue Yang4, Kaixin Ma2, Xiaoman Pan2, Yangqiu Song1, and Dong Yu2 1CSE Department, HKUST, 2Tencent AI Lab, Bellevue, USA 3Robotics X, Tencent, 4University of Pennsylvania zwanggy@cse.ust.hk, hongmzhang@global.tencent.com
@@ 1 Introduction
Large Vision-Language Models (LVLMs), such as Qwen2.5-VL Bai et al. (2025), GPT-4o OpenAI (2023), and others Zhu et al. (2025), have demonstrated state-of-the-art performance on a variety of vision-and-language tasks. However, their ability to comprehend embodied environments and navigate within them has been less explored, primarily due to the limited diversity of scenes and objects in current navigation benchmarks. For example, Matterport-3D (Chang et al., 2017) only considers 21 types of target objects in 90 private homes, and ProcTHOR (Deitke et al., 2022) contains 16 object types in four kinds of rooms (i.e., bedroom, living room, kitchen, and bathroom).
Figure 1: Illustration of our open-vocabulary object navigation. When fine-tuning LVLMs, we build CoT explanation traces to help them better grasp the rationale of object navigation.
@@
In this work, we propose the new task of open-vocab object navigation and first study LVLMs’ performance on the task, where an agent is required to navigate to a wide range of target objects without a pre-defined vocabulary. With this purpose, we introduce a new dataset, DivScene, which features the most comprehensive range of scene types and target objects to the best of our knowledge. Specifically, we collect 81 scene types based on the MIT Scenes Dataset (Quattoni and Torralba, 2009). Then, we use LLMs to automatically compose diverse house descriptions by adding attributes to those scene types, such as “a bakery with tile-patterned walls.” We input these descriptions into a language-guided framework, Holodeck (Yang et al., 2024), to build houses automatically with the strong ability of GPT-4 (OpenAI, 2023). In total, we compile 4,614 houses across 81 distinct scene types on the AI2THOR platform (Kolve et al., 2017). For benchmarking and training LVLMs, we further sampled shortest-path episodes in the houses from DivScene. Specifically, we discretize houses into grid maps with a fixed step size and randomly sample target objects in each house. Then, we search for the shortest paths from the agent’s initial position to target objects with BFS. In total, we collect about 23K shortest-path episodes, forming the DivSceneep. In total, our episode data contains 5,707 types of target object, significantly surpassing existing object navigation datasets Deitke et al. (2020); Chang et al. (2017). This makes it a comprehensive testbed for evaluating the open-vocab navigation capabilities.
With our collected data, we benchmarked various methods based on LVLMs and LLMs, including blind LLMs, LLMs with image captioning, open-source LVLMs, and proprietary LVLMs. Our results show that most current models are still unable to explore environments and navigate toward open-vocabulary objects. Specifically, most models failed to outperform a random baseline, while GPT-4o achieves a success rate of slightly above 30%, far from adequate for real-world applications. To enhance their navigation ability, we further studied fine-tuning LVLMs on the data sampled in our DivScene. Specifically, we propose a new navigational LVLM, called NatVLM (Navigational Chain-of-Thought VLM), which is fine-tuned from Idefics 2 (Laurençon et al., 2024) on our sampled episodes in DivSceneep. As shown in Figure˜1, the model is tuned to process the current observation, such as the egocentric view and the agent’s status, and generate the next action. We train Idefics 2 with imitation learning (IL; Brohan et al., 2022) using BFS-generated shortest paths in DivSceneep. To further improve the navigation ability, we also manually collect CoT explanation traces of each action prediction (Mitra et al., 2023; Ho et al., 2023) to help Idefics 2 understand the underlying rationale behind the navigation task.
In our experiments, we surprisingly discovered that simply imitating the shortest paths constructed by breadth-first search can be an effective approach to enhance the open-vocab navigation ability of LVLMs. Our NatVLM model outperforms off-the-shelf LVLMs and LLMs by a large margin, achieving a success rate approximately 20% higher than GPT-4o. Compared to existing IL-based methods, which rely on training models with large corpora of costly human demonstrations (Brohan et al., 2022; Wei et al., 2023), our approach offers a far more efficient and economical alternative. Further, we carry out thorough ablation studies to show the efficacy of CoT explanation traces in action prediction. Moreover, few-shot experiments demonstrate the robustness of our agent. Last but not least, we validate the generalization ability of our agent on three unseen datasets: ProcTHOR (Deitke et al., 2022), iTHOR (Weihs et al., 2021), and HM3D Ramakrishnan et al. (2021).
@@ 2 Related Work
Object Navigation: The Reinforcement Learning has long been used for tackling object navigation tasks (Zhu et al., 2017; Druon et al., 2020; Ehsani et al., 2021). However, recent studies (Ehsani et al., 2023) show that RL requires extensive reward shaping, limiting its practicality. Alternatively, imitation learning (Pomerleau, 1988; Zhang et al., 2018) has also been applied to these tasks, inspiring many subsequent works (Brohan et al., 2022, 2023; Ehsani et al., 2023). However, existing works only focus on closed-vocab object navigation tasks, mainly due to the limited diversity of current benchmarks. For example, Matterport-3D (Chang et al., 2017) only considers 21 kinds of target objects in private homes. ProcTHOR (Deitke et al., 2023a) solely contains 16 target objects in four kinds of rooms. Although HM3D-OVON (Yokoyama et al., 2024b) attempts to broaden the diversity, the number of object categories is still limited to 379, and there are only 181 scenes, making it difficult to fully reflect the open-vocab nature of real-world scenarios. In constrast, we introduce DivScene, a dataset comprising 5,707 categories of target objects among 22,696 object types, spanning 4,614 distinct scenes. Therefore, our dataset reflects a more diverse nature of real-world settings with an open vocabulary.
@@
VLMs and LLMs for Embodiment: LLMs and VLMs (Achiam et al., 2023; Touvron et al., 2023; Lu et al., 2024) have emerged as the foundation for solving embodied tasks (Pan et al., 2023; Majumdar et al., 2024). A few methods have employed contrastive VLMs (Radford et al., 2021; Li et al., 2022) as the visual encoders (Khandelwal et al., 2022; Majumdar et al., 2022) or object-grounding tools (Gadre et al., 2023; Dorbala et al., 2023) for navigation. Yu et al. (2023) utilizes an LLM as the planning backbone with a captioning model for perception. Following them, many improvements have been proposed (Cai et al., 2024; Chen et al., 2023; Zhou et al., 2023; Shah et al., 2023). While these works still focus on a small set of target objects, we explore the potential of recent LVLMs for open-vocabulary object navigation.
Figure 2: Data collection process. On the left, we show the process of collecting scenes. We prompt GPT-4o to collect textual house descriptions and use Holodeck to build houses. On the right, we show an episode built in the house, where we use BFS to find the shortest path. Then, actions and observations are collected.
@@ 3 Task Definition
We study the open-vocab object navigation task involving an agent in an environment to find the target object belonging to a given category. The object categories in the open-vocabulary setting are denoted by $C=\{c_{0},c_{1},c_{2},\ldots\}$, which is not restricted to a predefined set. The houses can be described by $H=\{h_{0},h_{1},\ldots,h_{n}\}$, and $n$ is the total number of scenes. In each episode, the embodied agent is initialized at a random position $p_{i}$ with rotation $r_{i}$ in a house $h_{i}$. Then, the embodied agent is required to perform instance-level object navigation, where it gets a target object within the category $c_{i}$ at the position $o_{i}$. Thus, an episode can be represented as $E^{(i)}=\{h_{i},p_{i},r_{i},c_{i},o_{i}\}$. At each time step $t$, the embodied agent observes the environment and predicts the next action $a_{t}$. Following previous work (Yu et al., 2023; Zhu et al., 2024), the observation comprises an RGB image of the egocentric view and the agent status (i.e., its position and rotation).
Our new houses are all collected on the AI2THOR platform (Kolve et al., 2017), and thus the action space, signified as $A$, covers: MoveAhead, RotateRight, RotateLeft, and Done. By default, the MoveAhead action moves the agent 25 centimeters, and RotateLeft (or RotateRight) rotates the agent 90 degrees. The Done action is used by the agent to indicate that the navigation task is complete. When the agent takes the Done action or reaches the max action limit, the episode is considered successful if the distance to the target object is below 1.5 meters. Based on the actions, an environment can be discretized as a 0.25 $\times$ 0.25-meter grid map of all reachable positions.
4 Data Collection for DivScene and DivSceneep
Existing object navigation studies only focus on limited types of scenes and objects. To fill this gap, we first curate a large-scale scene dataset DivScene, featuring 4,614 scenes. Then, 23K shortest-path episodes with 5,707 kinds of target objects were sampled using breadth-first search (BFS), forming the DivSceneep dataset. We illustrate the details in Figure˜2.
@@ 4.1 Scene Collection for DivScene
We adopt Holodeck (Yang et al., 2024) to build scenes, easing human labor. Holodeck takes textual house descriptions as input and uses GPT-4 to decide the layout, styles, and object selections. To collect diverse houses, we first manually compile 81 scene types across five categories by supplementing the MIT scene dataset (Quattoni and Torralba, 2009), like music studio and home office.
Figure 3: Examples of houses with different scene types.
@@
Then, we build textual house descriptions based on randomly chosen scene types by adding house attributes. We consider 12 house attributes, such as room style, users of the room, etc. We randomly sample 1-3 attributes and prompt GPT-4 to assign specific values to them. Given the scene type and attribute values, a house description is then written by GPT-4. Here, we prompt GPT-4 under the in-context learning setting (Brown, 2020), and five exemplars are provided to generate the description for the final test example. See full details of scene types and attributes in Section˜A.1 and concrete prompts in Section˜A.2. With a strict data filtering (Section˜A.3), we obtained 4,614 houses. We present a few houses in Figure˜3 and more examples in Appendix˜E.
4.2 Episode Collection for DivSceneep
@@
There are three steps in our episode sampling method. First, we sample the agent’s initial position and the target object. Then, we use BFS to find the shortest paths in the 0.25 $\times$ 0.25-meter grid map. Finally, we obtain the action sequence and corresponding observations. We show an example on the right side of Figure˜2.
@@
First, as discussed in Section˜3, houses are discretized as grid maps of the fixed step size. We randomly sample an initial position from the grip map for the agent. Then, a target object is randomly sampled from all available objects in the environment. To encourage diversity, objects of the same type as those sampled in previous episodes are removed from the pool when sampling for new episodes in the same house. We also impose that the target objects are between 0.3m and 2.0m in height to ensure they are observable to the agent.
@@
Second, we find the shortest path in the grid map that navigates a multi-room, cluttered environment. We can obtain the ground truth information from AI2-THOR, like reachable positions and objects’ coordinates. We use BFS to find the shortest path from the initial position to the target object. More details are in Section˜A.4. This ground truth information is not provided to agents for inference and is only used to produce episodes for training.
@@
With the shortest path, we then derive the sequence of actions needed to achieve navigation. Basically, between two adjacent positions in the shortest path, we add a MoveAhead action if the agent’s rotation remains unaltered. Otherwise, we first use RotateRight or RotateLeft to adjust the orientation and then add a MoveAhead action (more details in Section˜A.5). Thus, we obtain an action sequence that can steer the agent to the target object. We execute all actions in AI2THOR and collect observations at each step, including the agent’s status and egocentric images.
Metric ProcTHOR iTHOR DivScene
Scene types 4 4 81
Rooms per house 3.78 1.00 2.35
Room size (m2) 25.21 34.24 30.15
Obj. types per house 16.25 30.92 32.17
Obj. per house 35.64 47.26 111.42
Obj. types 38 116 22,696
Table 1: A comparison of ProcTHOR, iTHOR, and our dataset. “Obj.” refers to objects.
4.3 Statistics and Comparison